# Trading in CEXs and DEXs with Priority Fees and Stochastic Delays

### v1

@@ Abstract
We develop a mixed control framework that combines absolutely continuous controls with impulse interventions subject to stochastic execution delays. The model extends current impulse control formulations by allowing (i) the controller to choose the mean of the stochastic delay of their impulses, and allowing (ii) for multiple pending orders, so that several impulses can be submitted and executed asynchronously at random times. The framework is motivated by an optimal trading problem between centralized (CEX) and decentralized (DEX) exchanges. In DEXs, traders control the distribution of the execution delay through the priority fee paid, introducing a fundamental trade-off between delays, uncertainty, and costs. We study the optimal trading problem of a trader exploiting trading signals in CEXs and DEXs. From a mathematical perspective, we derive the associated dynamic programming principle of this new class of impulse control problems, and establish the viscosity properties of the corresponding quasi-variational inequalities. From a financial perspective, our model provides insights on how to carry out execution across CEXs and DEXs, highlighting how traders manage latency risk optimally through priority fee selection. We show that employing the optimal priority fee has a significant outperformance over non-strategic fee selection.
@@ 1 Introduction
The interplay between trading speed, execution uncertainty, and market structure has become increasingly prominent with the rise of decentralized exchanges (DEXs). Currently, within the crypto space, DEXs operate alongside traditional centralized exchanges (CEXs). In DEXs, traders face a fundamental trade-off between immediacy and execution risk, as the timing (position in queue) of order execution is subject to a stochastic delay that can be shortened by offering a “priority fee” to the miners. This motivates the need for the development of mathematical tools capable of capturing decision-making under delayed (and to some extent controlled) uncertain executions. In this paper, we develop theory for a new type of impulse control problem with stochastic delay in which the controller can influence the stochastic delay of the execution of their impulses through a control variable. We extend the literature in two key directions: (i) enabling control over the mean of the stochastic delay of the impulses, and (ii) allowing for multiple pending orders within the impulse control with stochastic delay framework. From a financial perspective, our framework provides a novel formulation for optimal trading between CEXs and DEXs, where the trader strategically selects priority fees to manage execution speed and risk.
@@
A detailed summary of our contributions is as follows. Mathematically, we introduce a framework that handles multiple pending orders within the class of impulse control problems with stochastic delay. Furthermore, we allow the controller to choose the mean of the random delay of each of their impulses. We combine a regular control, which models continuous trading on the CEX, with an impulse control, which models discrete order submissions to the DEX. Technically, adding the regular control means that between two impulse decisions, the state is continuously controlled with its own associated optimization. On the impulse side, closest to our work are [20, 19, 7], where the authors investigate optimal stopping/impulse control problems with deterministic delay. In [19], an infinite-horizon setting is considered with deterministic delay and an arbitrary number of pending orders, while [7] treats a finite-horizon problem with any finite number of pending orders. In these deterministic-delay models, the main technical burden comes from enforcing the associated boundary and consistency conditions induced by the fixed delay. By contrast, [11] consider stochastic execution delay where the delay mechanism is exogenous and not a control variable. The controller chooses order submission times and sizes, while execution latency is governed by a given Poisson process, leading to a different structure in the intervention operator. The closest to our work is [11], where the authors introduced an impulse control with uncontrolled stochastic delay and one pending order. Here, on the other hand, we control the mean of the stochastic delay and we allow for a finite number of pending orders. Regarding the option of the controller to choose the mean of the stochastic delay, the closest work we are aware of is [3], who deal with a Markov decision process and a discrete-time version of a controlled deterministic delay (in their framework this is called observation delay). As far as we are aware, our framework is the first to address optimal execution with random latency in which the trader can influence the mean of the latency distribution.
@@
From a financial point of view and to the best of our knowledge, our work is the first to study optimal trading between centralised exchanges (CEX) and decentralised exchanges (DEX) employing the key degree of freedom that liquidity takers have when choosing the “priority fee” of their orders. We find the optimal decision boundaries that traders should employ when selecting the priority fee attached to their orders. Furthermore, we show how these regions change with time, inventory, and price discrepancies. As expected, the performance of the trader increases with the number of priority fees they consider, but we find that this performance plateaus fairly quickly, which implies that entertaining a finite number of priority fees is close to optimal. Within the CEX-DEX trading context, closest to our work are [6, 8, 9, 15, 17, 18]. However, none of these works accounted for execution delay and the crucial (and yet understudied) role of the priority fees. To the best of our knowledge, this is the first framework to study trading in DEXs accounting for priority fees and execution delay.
@@
The remainder of the paper proceeds as follows. Section 2 formulates the mixed control problem with stochastic execution delay. Section 3 establishes the dynamic programming principle and characterizes the value function via the associated Hamilton–Jacobi–Bellman quasi-variational inequality, showing that it is the unique viscosity solution of the HJBQVI. Section 4 applies the framework to an optimal trading problem in which the agent trades continuously on a centralized exchange and discretely on a decentralized exchange. It also provides numerical illustrations and studies how the optimal strategy varies across model specifications and parameter choices.
@@ 2 Problem Formulation
The proof is standard and it follows closely the arguments in [22, Chapter V]. ∎
@@ 3.1 Dynamic programming principle
is progressively measurable. Feasibility and integrability hold on $t,\tau)$ by admissibility of $\alpha$, and on $\tau,T]$ by admissibility of $\bar{\alpha}^{\varepsilon}$. Additionally, the number of impulses is respected because $\bar{\alpha}^{\varepsilon}\in\mathcal{A}_{K,p(\tau,\alpha)}(\tau)$. The non-accumulation property is preserved by construction (see [19, Chapter IX]). Now using the law of iterated conditional expectations and (10), we get that
@@ 4.2 Numerical Results
Figure 3 explores the effect of the inventory in the priority fees. We take the inventory to be large and positive $q=40$ (negative case is analogous).
@@
Figure 3: Exercise and continuation regions as a function of the CEX and DEX prices $(s,z)$ at time indices $t_{0}<t_{1}<t_{2}$ for $q=40$. Here, $t_{0}=0.2$, $t_{1}=0.5$, and $t_{2}=0.8$.
@@ 5 Conclusions
We introduced a new type of mixed control problem where the agent is allowed to choose the expected delay of the execution of the impulses. The associated value function to this optimization problem is shown to satisfy a system of variational Hamilton–Jacobi–Bellman inequalities in the viscosity sense. Moreover, we establish uniqueness of the solution to this HJB-QVI. We apply our formulation to the problem of optimal trading between CEX and DEX, where the agent chooses the priority fee attached to the orders sent to the DEX. We find the optimal priority fee as a function of time, inventories, and price dislocations between the CEX and the DEX. Our results show that the outperformance one gets from employing more priority fees rapidly plateaus, and that the optimal fee selection brings about significant outperformance over a non-strategic fee selection. To the best of our knowledge, this is the first time this appears in the growing literature on AMMs.
@@ A.1 Viscosity Solution
$\begin{aligned} &\frac{\gamma_{m}}{h_{m}}+\mathbb{E}\Bigg[\frac{1}{h_{m}}\int_{t_{m}}^{\tilde{t}_{m}}\bigg(\Big(-\frac{\partial\varphi}{\partial t}-\mathcal{L}^{a}\varphi\Big)(s,X^{t_{m},x_{m},\alpha_{m}}_{s},p(s,\alpha_{m}))-f\big(s,X^{t_{m},x_{m},\alpha_{m}}_{s},a\big)+\sum_{i=1}^{N}\mathds{1}_{\{\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle>0\}}\,\ell_{i}c\big(s,X^{t_{m},x_{m},\alpha_{m}}_{s},\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle,\langle\mathfrak{v}(s,\alpha_{m}),e_{i}\rangle\big)\\ &-\sum_{i=1}^{N}\mathds{1}_{\{\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle>0\}}\,\ell_{i}\Big(\varphi\big(s,\Gamma(s,X^{t_{m},x_{m},\alpha_{m}}_{s},\langle\mathfrak{v}(s,\alpha_{m}),e_{i}\rangle),\mathfrak{i}(s,\alpha_{m})-\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle e_{i},\mathfrak{v}(s,\alpha_{m})-\langle\mathfrak{v}(s,\alpha_{m}),e_{i}\rangle e_{i}\big)-\varphi\big(s,X^{t_{m},x_{m},\alpha_{m}}_{s},p(s,\alpha_{m})\big)\Big)\bigg)\mathrm{d}s\,\Bigg]\geq 0,\end{aligned}$
@@ A.3 Uniqueness and Continuity Result
Moreover, since the above inequality remains valid when the supremum in the Hamiltonian is restricted to a sufficiently large compact set, the sequence $(a_{k}^{\epsilon})_{k\in\mathbb{N}}$ may be chosen bounded. By the Bolzano-Weierstrass theorem, there exists a subsequence, still denoted $(a_{k}^{\epsilon})_{k\in\mathbb{N}}$, converging to some $a^{\epsilon}_{0}\in\mathbb{R}$. Sending $k\to+\infty$ is therefore justified for fixed $R$. Letting $R\to+\infty$ and using the monotone convergence $H_{R}\underset{R\to+\infty}{\to}H$ (see [14, Chapter III]), we recover the inequality for the original Hamiltonian. Sending $k$ to $\infty$, we get by continuity of $b$ and $\sigma$ and inequality (35) that

### v2

@@ Abstract
We develop a mixed control framework that combines absolutely continuous controls with impulse interventions subject to stochastic execution delays. The model extends current impulse control formulations by allowing (i) the controller to choose the mean of the stochastic delay of their impulses, and allowing (ii) for multiple pending orders, so that several impulses can be submitted and executed asynchronously at random times. The framework is motivated by an optimal trading problem between centralized (CEX) and decentralized (DEX) exchanges. In DEXs, traders control the distribution of the execution delay through the priority fee paid, introducing a fundamental trade-off between delays, uncertainty, and costs. We study the optimal trading problem of an agent exploiting trading signals in CEXs and DEXs. From a mathematical perspective, we derive the associated dynamic programming principle of this new class of impulse control problems, and establish the viscosity properties of the corresponding quasi-variational inequalities. From a financial perspective, our model provides insights on how to carry out execution across CEXs and DEXs, highlighting how traders manage latency risk optimally through priority fee selection. We show that employing the optimal priority fee has a significant outperformance over non-strategic fee selection.
@@ 1 Introduction
The interplay between trading speed, execution uncertainty, and market structure has become increasingly prominent with the rise of decentralized exchanges (DEXs). Currently, within the crypto space, DEXs operate alongside traditional centralized exchanges (CEXs). In DEXs, traders face a fundamental trade-off between immediacy and execution risk, as the timing (position in queue) of order execution is subject to a stochastic delay that can be shortened by offering a “priority fee” to the miners. This motivates the need for the development of mathematical tools capable of capturing decision-making under delayed (and to some extent controlled) uncertain executions. In this paper, we develop a theory for a new type of impulse control problem with stochastic delay in which the controller influences the stochastic delay of the execution of their impulses through a control variable. We extend the literature in two key directions: (i) enabling control over the mean of the stochastic delay of the impulses, and (ii) allowing for multiple pending orders within the impulse control with stochastic delay framework. From a financial perspective, our framework provides a novel formulation for optimal trading between CEXs and DEXs, where the trader strategically selects priority fees to manage execution speed and risk.
@@
A detailed summary of our contributions is as follows. Mathematically, we introduce a framework that handles multiple pending orders within the class of impulse control problems with stochastic delay. Furthermore, we allow the controller to choose the mean of the random delay of each of their impulses. We combine a regular control, which models continuous trading on the CEX, with an impulse control, which models discrete order submissions to the DEX. On the impulse side, [24, 23, 7, 14], investigate optimal stopping/impulse control problems with deterministic delay. In [23], an infinite-horizon setting is considered with deterministic delay and an arbitrary number of pending orders, while [7] treats a finite-horizon problem with any finite number of pending orders. In these deterministic-delay models, the main technical burden comes from enforcing the associated boundary and consistency conditions induced by the fixed delay. [^1] By contrast, [14] consider stochastic execution delay where the delay mechanism is exogenous and not a control variable. The controller chooses order submission times and sizes, while execution latency is governed by a given Poisson process, leading to a different structure in the intervention operator. Closest to our work is [14], where the authors introduced an impulse control problem with uncontrolled stochastic delays and one pending order. Here, on the other hand, we control the mean of the stochastic delay and we allow for a finite number of pending orders. Regarding the option of the controller to choose the mean of the stochastic delay, the only work that we are aware of is [3], who deal with a Markov decision process and a discrete-time version of a controlled deterministic delay (in their framework this is called observation delay). As far as we are aware, our framework is the first to address optimal execution with random latency in which the trader can influence the mean of the latency distribution.
@@
From a financial point of view and to the best of our knowledge, our work is the first to study optimal trading between centralised exchanges (CEX) and decentralised exchanges (DEX) while employing the key degree of freedom that liquidity takers have when choosing the “priority fee” of their orders. We find the optimal decision boundaries that traders should employ when selecting the priority fee attached to their orders. Furthermore, we show how these regions change with time, inventory, and price discrepancies. As expected, the performance of the trader increases with the number of priority fees they consider, but we find that this performance plateaus fairly quickly, which implies that entertaining a finite number of priority fees is close to optimal. Within the CEX-DEX trading context, closest to our work are [6, 11, 12, 18, 21, 22]. However, none of these works accounted for execution delays and the crucial role of the priority fees. To the best of our knowledge, this is the first continuous-time optimal control framework to tackle the problem of trading in CEX-DEXs accounting for both priority fees and execution delays. [^2] Priority fees have been studied in [10, 9] within the context of price discovery and market efficiency. More precisely, [10] introduce a competition model for trading in blockchains with priority fees and its impact on price discovery, and [9] investigates whether priority fees carry information (or market impact). On the other hand, [20, 1, 2, 8] study the role of fees in attracting order flow to the venue.
@@
The remainder of the paper proceeds as follows. Section 2 formulates the mixed control problem with (controlled) stochastic execution delay. Section 3 establishes the dynamic programming principle and characterizes the value function via the associated Hamilton–Jacobi–Bellman quasi-variational inequality, showing that it is the unique viscosity solution of the HJBQVI. Section 4 applies the framework to an optimal trading problem in which the agent trades continuously on a centralized exchange and discretely on a decentralized exchange. Lastly, we provide numerical illustrations and study how the optimal strategy varies across model specifications and parameter choices.
@@ 2 Problem Formulation
The proof is standard and it follows closely the arguments in [26, Chapter V]. ∎
@@ 3.1 Dynamic programming principle
is progressively measurable. Feasibility and integrability hold on $t,\tau)$ by admissibility of $\alpha$, and on $\tau,T]$ by admissibility of $\bar{\alpha}^{\varepsilon}$. Additionally, the number of impulses is respected because $\bar{\alpha}^{\varepsilon}\in\mathcal{A}_{K,p(\tau,\alpha)}(\tau)$. The non-accumulation property is preserved by construction (see [23, Chapter IX]). Now using the law of iterated conditional expectations and (10), we get that
@@ 4.2 Numerical Results
Figure 3 explores the effect of the inventory in the priority fees. We take the inventory to be large and positive $q=20$ (negative case is analogous).
@@
Figure 3: Exercise and continuation regions as a function of the CEX and DEX prices $(s,z)$ at time indices $t_{0}<t_{1}<t_{2}$ for $q=20$. Here, $t_{0}=0.2$, $t_{1}=0.5$, and $t_{2}=0.8$.
@@ 5 Conclusions
We introduced a new type of mixed control problem where the agent is allowed to choose the expected delay of the execution of their impulses. The associated value function to this optimization problem is shown to satisfy a system of variational Hamilton–Jacobi–Bellman inequalities in the viscosity sense. Moreover, we establish uniqueness of the solution to this HJB-QVI. We apply our formulation to the problem of optimal trading between CEX and DEX, where the agent chooses the priority fee attached to the orders sent to the DEX. We find the optimal priority fee as a function of time, inventories, and price dislocations between the CEX and the DEX. Our results show that the outperformance one gets from employing more priority fees rapidly plateaus, and that the optimal fee selection brings about significant outperformance over a non-strategic fee selection.
@@ A.1 Viscosity Solution
$\begin{aligned} &\frac{\gamma_{m}}{h_{m}}+\mathbb{E}\Bigg[\frac{1}{h_{m}}\int_{t_{m}}^{\tilde{t}_{m}}\bigg(\Big(-\frac{\partial\varphi}{\partial t}-\mathcal{L}^{a}\varphi\Big)(s,X^{t_{m},x_{m},\alpha_{m}}_{s},p(s,\alpha_{m}))-f\big(s,X^{t_{m},x_{m},\alpha_{m}}_{s},a\big)+\sum_{i=1}^{N}\mathds{1}_{\{\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle>0\}}\,\ell_{i}c\big(s,X^{t_{m},x_{m},\alpha_{m}}_{s},\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle,\langle\mathfrak{v}(s,\alpha_{m}),e_{i}\rangle\big)\\ &-\sum_{i=1}^{N}\mathds{1}_{\{\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle>0\}}\,\ell_{i}\Big(\varphi\big(s,\Gamma(s,X^{t_{m},x_{m},\alpha_{m}}_{s},\langle\mathfrak{v}(s,\alpha_{m}),e_{i}\rangle),\mathfrak{i}(s,\alpha_{m})-\langle\mathfrak{i}(s,\alpha_{m}),e_{i}\rangle e_{i},\mathfrak{v}(s,\alpha_{m})-\langle\mathfrak{v}(s,\alpha_{m}),e_{i}\rangle e_{i}\big)-\varphi\big(s,X^{t_{m},x_{m},\alpha_{m}}_{s},p(s,\alpha_{m})\big)\Big)\bigg)\mathrm{d}s\,\Bigg]\\ &\geq 0,\end{aligned}$
@@ A.3 Uniqueness and Continuity Result
Moreover, since the above inequality remains valid when the supremum in the Hamiltonian is restricted to a sufficiently large compact set, the sequence $(a_{k}^{\epsilon})_{k\in\mathbb{N}}$ may be chosen bounded. By the Bolzano-Weierstrass theorem, there exists a subsequence, still denoted $(a_{k}^{\epsilon})_{k\in\mathbb{N}}$, converging to some $a^{\epsilon}_{0}\in\mathbb{R}$. Sending $k\to+\infty$ is therefore justified for fixed $R$. Letting $R\to+\infty$ and using the monotone convergence $H_{R}\underset{R\to+\infty}{\to}H$ (see [17, Chapter III]), we recover the inequality for the original Hamiltonian. Sending $k$ to $\infty$, we get by continuity of $b$ and $\sigma$ and inequality (35) that